Calculate the vapor pressure of a solution at 35°C containing 250-g of a nonelectrolyte solute (FM = 350 g/mol) in 1750-mL of the solvent. The vapor pressure of the pure solvent at this temperature is 50 mm Hg. The density of the pure solvent is 3.5 g/mL and its molar mass is 40 g/mol. 



1
Expert's answer
2021-02-09T04:38:47-0500

Vapor pressure of pure solvent(P°) = 50 mm Hg

Mass of solvent = 1750 × 3.5

= 6125 g

=

Moles of solute = 250/350 = 0.714 mol

Moles of solvent = 6125/40 = 153.1 mol

Using Vapor pressure formula

"(P\u00b0-P)\/P\u00b0 = X"

Where, X is mole fraction of solute

(50 - P)/50 = 0.714/(0.714+153.1)

P = 49.8 torr

Vapor pressure of solution = [49.8] torr